OpenStudy (anonymous):

Evaluate the function f(x) at the given numbers (correct to six decimal places). f(x) = x^2 − 3x/x^2 − x − 6 , x = 3.5

OpenStudy (anonymous):

is it 0.636364

jhonyy9 (jhonyy9):

f(x)= (x(x-3))/((x-3)(x+2)) = can you continue it ?

OpenStudy (anonymous):

you get the same answer

jhonyy9 (jhonyy9):

continue what i have started

OpenStudy (anonymous):

don't need to simplify

jhonyy9 (jhonyy9):

by simplifie and will get easy the right result

OpenStudy (anonymous):

i'm just confused about rounding my answer

jhonyy9 (jhonyy9):

so simplifie by x-3 and see than what will get

OpenStudy (anonymous):

okaaay

OpenStudy (anonymous):

same answer

OpenStudy (anonymous):

0.636363636

jhonyy9 (jhonyy9):

so than will get x/(x+2) = if x=3,5 than will get 3,5/(3,5 +2) =3,5 / 5,5 =

jhonyy9 (jhonyy9):

=0,636363... yes you are right

OpenStudy (anonymous):

do i round it up

OpenStudy (anonymous):

0.636364

jhonyy9 (jhonyy9):

ok good luck bye